ssddss02:

What is 345x times -332y. When x = 8 and y = 5?

wo1f0mon:

so 345*8-332*5

wo1f0mon:

2760-1660

ssddss02:

Ok.

wo1f0mon:

1100

KjSaif:

^^

ssddss02:

Ok, I need work to be shown aswell.

wo1f0mon:

ok wait

ssddss02:

Ok, waiting.

KjSaif:

you deserve a star/medal xD @wo1f0mon

wo1f0mon:

345*8-332*5 =2760-1660 =1100

ssddss02:

No, full work showing everything.

wo1f0mon:

ok

ssddss02:

yay!! I'm 50 SS

Ultrilliam:

I got dis.

KjSaif:

congrats xD @ssddss02

ssddss02:

Thanks!

ssddss02:

Now 51!

wo1f0mon:

waIT

ssddss02:

Ok, waiting.

Ultrilliam:

when you see that variable after the number, it means you should multiply the number by the variable.

wo1f0mon:

345x * (-332y)

ssddss02:

ok

wo1f0mon:

CHILL AND WAIT FOR ME

ssddss02:

I am I am!!

KjSaif:

xD

Ultrilliam:

so that means the first step is to get 345*x and since x is 8 that means 345*8, so what is 345 times 8?

ssddss02:

Ok ultrill .

wo1f0mon:

2760 +(-1660) =1100

ssddss02:

wolf ultrills beating you to it.

wo1f0mon:

@Ultrilliam I FINISHED

ssddss02:

Ok.

Ultrilliam:

the difference is, you just gave a direct answer, where as I am explaining how to solve it for future problems.

ssddss02:

Thanks.

wo1f0mon:

OH SORRY

ssddss02:

I have to show work If I didn't do the problem how do I know how to show the work?

wo1f0mon:

wot happen again plz dont ban me

KjSaif:

thanks @Ultrilliam and @wo1f0mon for helping @ssddss02 even though @KjSaif could have helped

Ultrilliam:

LOL

KjSaif:

xD

wo1f0mon:

XD

ssddss02:

Yes you could of but you chose to spam and not to!

ssddss02:

I think I understand it now thanks.

wo1f0mon:

good luck with the other qustions

ssddss02:

Thanks.